Phillips Company manufactures air-conditioning units for commercial buildings and has noticed considerable variation in shipping expenses from month to month as per the data below:

Units Total Shipping
  Month Shipped Expense
  January 4 $ 2,280
  February 4 $ 3,180
  March 5 $ 2,700
  April 2 $ 1,302
  May 3 $ 2,250
  June 6 $ 3,120
  July 8 $ 4,512

4. What is the cost formula for shipping expenses?

Y =    + X

5. If seven air conditioners are shipped in August, estimate the total shipping expenses.

Total shipping expense

6. If nine air conditioners are shipped in September, estimate the fixed shipping expenses.

Fixed expense

7. If 10 air conditioners are shipped in October, estimate the variable shipping expense.

Variable expense

8. If the air conditioners have an average sales price of $7,500, variable manufacturing costs are $3,750 per unit, variable manufacturing overhead is $1,500 per unit, and variable selling and administration costs (excluding shipping) are $300 per unit, what is the contribution margin per unit?

Contribution margin per unit

9. Given the facts in part 3–8, estimate the total contribution margin for August if seven air conditioners are produced and sold.

Total contribution margin

10. Given the facts in part 3–8, if total fixed costs (excluding shipping) are $1,850 per month, estimate operating income for August if seven air conditioners are produced and sold.

Operating income
Answer #1

  Month Units shipped (x) Expense (y) x*y x*x y*y
  January 4 2,280              9,120                   16           5,198,400
  February 4 3,180           12,720                   16         10,112,400
  March 5 2,700           13,500                   25           7,290,000
  April 2 1,302              2,604                     4           1,695,204
  May 3 2,250              6,750                     9           5,062,500
  June 6 3,120           18,720                   36           9,734,400
  July 8 4,512           36,096                   64         20,358,144
32             19,344           99,510                 170         59,451,048
n 7
n*Sum(xy)           696,570
Sumx*Sumy           619,008
n*sumx2               1,190
(Sumx)2               1,024

b = (696570 - 619008)/(1190-1024) = 467.24

a = (19344 - 467.24*32)/7 = 627.47

4) y = 627.47 + 467.24*x

5) For x = 7, total expense y = 627.47 + 467.24*7 = 3898.16 $

6) The fixed expense = a = 627.47 $ (It will remain same irrespective of number of AC shipped)

7) The variable expense = b*x

for x = 10, Variable expense = 467.24*10 = 4672.4 $